Issue Price of a Bond
Abbott, Inc., plans to issue $500,000 of ten percent bonds that will pay interest semiannually and mature in five years. Assume that the effective interest rate is 12 percent per year compounded semiannually. Calculate the selling price of the bonds.

Round answers to the nearest whole number.

Selling price of bonds is $
Answer #1
Amount PV factor 6% Present value
Semi-annual interest 25000 7.36009 184002.18
Principal 500000 0.55839 279197.39
Total 463199.56
Selling price of bonds is 463200 or 463199
Workings:
Semi-annual interest 25000 =500000*10%/2
Semi-annual effective rate 6% =12%/2
Semi-annual periods 10 =5*2
PV factor 6%:
Semi-annual interest 7.36009 =(1-(1.06)^-10)/0.06
Principal 0.55839 =1/1.06^10